OpenStudy (anonymous):

Luke has a spinner divided into ten equal segments, each with a different color. In fifty spins, he spun yellow ten times. How does this compare to the expected result? The experiment resulted in 5 more yellows than expected. The experiment resulted in ten more yellows than expected. The experiment resulted in 5 fewer yellows than expected. The experiment resulted in ten fewer yellows than expected.

OpenStudy (anonymous):

plz help me somebody

OpenStudy (whpalmer4):

If you have 10 equal segments on the spinner, and only 1 of them is yellow, what is the probability of getting yellow on 1 spin?

OpenStudy (anonymous):

1

OpenStudy (whpalmer4):

a probability of 1 means that it is certain: all the probabilities add to 1. Next guess? :-)

OpenStudy (anonymous):

3

OpenStudy (whpalmer4):

come on, if 1 is certain, what does 3 represent?

OpenStudy (anonymous):

it means un certain......I don't know im bad at math

OpenStudy (whpalmer4):

if you have a multiple choice question with 2 answers, both equally likely, and you guess, what is the probability of getting a correct answer?

OpenStudy (anonymous):

50%

OpenStudy (whpalmer4):

or flipping a coin: if I flip a coin right now, and you guess heads or tails, what is the probability that you will be right?

OpenStudy (whpalmer4):

okay, 50%. why 50%?

OpenStudy (anonymous):

because theirs only to so you would put it on half

OpenStudy (whpalmer4):

okay. there's 1 favorable outcome, but 2 possible outcomes. The probability of a favorable outcome is # of favorable outcomes / # of possible outcomes 1/2 = 0.5 or 50% now, look at that spinner problem again. What do you think the answer should be?

OpenStudy (whpalmer4):

(for what is the probability of getting yellow with 1 spin)

OpenStudy (anonymous):

the second one?

OpenStudy (anonymous):

no...........? am i wrong

OpenStudy (anonymous):

are you still their

OpenStudy (whpalmer4):

how many possible outcomes are there?

OpenStudy (whpalmer4):

10 different colors, each segment the same size...

OpenStudy (anonymous):

The answer would be the first one right

OpenStudy (whpalmer4):

by the way, as you've made this mistake a couple of times now, "their" is a possessive, like "those are their homework problems" to ask if someone or something is still present, "there" is the word you want. "Hello, are you there?" if you are saying "they are", then you can use "they're", which is a contraction.

OpenStudy (whpalmer4):

No, we're going to work through this, and you won't need to ask me which answer is correct, you'll understand and know which is correct :-)

OpenStudy (anonymous):

I know but I have to have the answer right now im in a rush

OpenStudy (whpalmer4):

so, we take 1 spin on that spinner. what is the probability that we get a yellow?

OpenStudy (whpalmer4):

colors are red, blue, green,yellow, orange, purple, magenta, violet, black, white only 1 is yellow, right? so we have a 1 in 10 chance. on average, if we spin it a large number of times, we should expect about 1/10 of the number to be yellow.

OpenStudy (whpalmer4):

so, let's say that 50 is a large enough number. if we spin 50 times, and we expect about 1/10 of the spins to be yellow, how many yellows is that?

OpenStudy (whpalmer4):

50 * 1/10 = 50/10 = 5, right? but we got 10 yellows, according to the problem statement, so that is 10 -5 = 5 more than we expected.

OpenStudy (anonymous):

oh i get it

OpenStudy (anonymous):

thanks for the help

OpenStudy (whpalmer4):

I appreciate that you're in a rush to get an answer, but OpenStudy isn't here to give you answers, it's here to help you learn.

OpenStudy (anonymous):

yep

OpenStudy (whpalmer4):

I'm happy to answer questions, but you'll have to work along with me, so be sure to allow plenty of time next time :-)